# Compositional Risk Minimization

### v1

@@ Abstract
In this work, we tackle a challenging and extreme form of subpopulation shift, which is termed compositional shift. Under compositional shifts, some combinations of attributes are totally absent from the training distribution but present in the test distribution. We model the data with flexible additive energy distributions, where each energy term represents an attribute, and derive a simple alternative to empirical risk minimization termed compositional risk minimization (CRM). We first train an additive energy classifier to predict the multiple attributes and then adjust this classifier to tackle compositional shifts. We provide an extensive theoretical analysis of CRM, where we show that our proposal extrapolates to special affine hulls of seen attribute combinations. Empirical evaluations on benchmark datasets confirms the improved robustness of CRM compared to other methods from the literature designed to tackle various forms of subpopulation shifts.
@@ 1 Introduction
We work with multi-attribute data, where each input (e.g., an image) is associated with multiple categorical attributes, and the task is to predict an attribute or multiple attributes. During training, we observe inputs from only a subset of all possible combinations of individual attributes, and during test we will see novel combinations of attributes never seen at training. Following, Liu et al. (2023), we refer to this distribution shift as compositional shift. These distribution shifts can also be viewed as an extreme case of subpopulation shift (Yang et al., 2023b). Towards the goal of tackling these compositional shifts, we develop an adaptation of naive discriminative Empirical Risk Minimization (ERM) tailored for multi-attribute data under compositional shifts. We term our approach Compositional Risk Minimization (CRM). The foundations of CRM are built on additive energy distributions that are studied in generative compositionality (Liu et al., 2022a), where each energy term represents one attribute. In CRM, we first train an additive energy classifier to predict all the attributes jointly, and then we adjust this classifier for compositional shifts.
@@
  • Theory of discriminative compositional shifts: For the family of additive energy distributions, we prove that additive energy classifiers generalize compositionally to novel combinations of attributes represented by a special mathematical object, which we call discrete affine hull. Our characterization of extrapolation is sharp, i.e., we show that it is not possible to generalize beyond discrete affine hull. We show that the volume of discrete affine hull grows fast in the number of training attribute combinations thus generalizing to many attribute combinations. The proofs developed in this work are quite different from existing works on distribution shifts and hence may be of independent interest.
@@
  • A practical method: CRM is a simple algorithm for training classifiers, which first trains an additive energy classifier and then adjusts the trained classifier for tackling compositional shifts. We empirically validate the superiority of the CRM algorithm to other algorithms previously proposed for robustness to various forms of subpopulation shifts.
2 Related Works
Compositional Generalization
Compositionality has long been seen as an essential capability (Fodor and Pylyshyn, 1988; Hinton, 1990; Plate et al., 1991; Montague, 1970) on the path to building human-level intelligence. The history of compositionality being too long to cover in detail here, we refer the reader to these surveys (Lin et al., 2023; Sinha et al., 2024). Most prior works have focused on generative aspect of compositionality, where the model needs to recombine individual distinct factors/concepts and generate the final output in the form of text (Gordon et al., 2019; Lake and Baroni, 2023) or image. For image generation in particular, a fruitful line of work is rooted in additive energy based models (Du et al., 2020, 2021; Liu et al., 2021; Nie et al., 2021), which translates naturally to additive diffusion models (Liu et al., 2022a; Su et al., 2024). Our present work also leverages an additive energy form, but our focus is on learning classifiers robust under compositional shifts, rather than generative models.
On the theoretical side, recently, there has been a growing interest in building provable approaches for compositional generalization (Dong and Ma, 2022; Wiedemer et al., 2023, 2024; Brady et al., 2023; Lachapelle et al., 2024). These works study models where the labeling function or the decoder is additive over individual features, and prove generalization guarantees over the Cartesian product of the support of individual features. While these works take promising and insightful first steps for provable compositional guarantees, the assumption of additive deterministic decoders may come as quite restrictive. In particular a given attribute combination can then only correspond to a unique observation, produced by a very limited interaction between generative factors, not to a rich distribution of observations. By contrast an additive energy model can associate an almost arbitrary distribution over observations to a given set of attributes. Based on this more realistic assumption of additive energy, our goal is to develop an approach that provably enables zero-shot compositional generalization in discriminative tasks, where the model needs to robustly predict never seen before factor combinations that the input is composed of.
Domain Generalization
Generalization under subpopulation shifts, where certain groups or combinations of attributes are underrepresented in the training data, is a well-known challenge in machine learning. Group Distributionally Robust Optimization (GroupDRO) (Sagawa et al., 2019) is a prominent method that minimizes the worst-case group loss to improve robustness across groups. Invariant Risk Minimization (IRM) (Arjovsky et al., 2019) encourages the model to learn invariant representations that perform well across multiple environments. Perhaps the simplest methods are SUBG and RWG (Idrissi et al., 2022), which focus on constructing a balanced subset or reweighting examples to minimize or eliminate spurious correlations. There are many other interesting approaches that were proposed, see the survey (Zhou et al., 2022) for details. The theoretical guarantees developed for these approaches (Rosenfeld et al., 2020; Arjovsky et al., 2019; Ahuja et al., 2020) require a large diversity in terms of the environments seen at the training time. In our setting, we incorporate inductive biases based on additive energy distributions that help us arrive at provable generalization with limited diversity in the environments.
Closely related to our proposed method are the logit adjustment methods Kang et al. (2019); Menon et al. (2020); Ren et al. (2020) used in robust classification. Kang et al. (2019) introduced Label-Distribution-Aware Margin (LDAM) loss for long-tail learning, proposing a method that adjusts the logits of a classifier based on the class frequencies in the training set to counteract bias towards majority classes. Closest to our work are the Logit Correction (LC) (Liu et al., 2022b) and Supervised Logit Adjustment (sLA) (Tsirigotis et al., 2024) methods that use logit adjustment for group robustness. LC adjusts logits based on the joint distribution of environment and class label, reducing reliance on spurious features in imbalanced training sets. Supervised Logit Adjustment (sLA) adjusts logits according to the conditional distribution of classes given the environment.
@@ 3 Problem Setting
3.1 Generalizing under Compositional Distribution Shift
@@
Formally, let $p(x,z)=p(z)p(x|z)$ denote the train distribution, and $q(x,z)=q(z)q(x|z)$ the test distribution. We denote the support of each attribute component $z_{i}$ under training distribution as $\mathcal{Z}_{i}^{\mathsf{train}}$ and the support of $z$ under training distribution as $\mathcal{Z}^{\mathsf{train}}$. The corresponding supports for the test distribution are denoted as $\mathcal{Z}_{i}^{\mathsf{test}}$ and $\mathcal{Z}^{\mathsf{test}}$. We define the Cartesian product of marginal support under training as ${\mathcal{Z}^{\times}}\coloneq\mathcal{Z}_{1}^{\mathsf{train}}\times\mathcal{Z}_{2}^{\mathsf{train}}\times\cdots\mathcal{Z}_{m}^{\mathsf{train}}$.
In this work, we study compositional shifts that are characterized by:
@@
The first point states that the conditional density of inputs conditioned on attributes remains invariant from train to test, which can be understood as the data generation mechanism from attributes to the inputs remains invariant. What changes between train and test is thus due to only shifting prior probabilities of attributes from $p(z)$ to $q(z)$. The second point specifies how these differ in their support: at test we observe novel combinations of individual attributes but not a completely new individual attribute. The task of compositional generalization is then to build classifiers that are robust to such compositional distribution shifts. Also, we remark that the above notion should remind the reader of the notion of Cartesian Product Extrapolation (CPE) from Lachapelle et al. (2024). Specifically, if a model succeeds on test distributions $q(z)$ that have a support equal to the full Cartesian product, i.e., $\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$, then it is said to achieve CPE.
@@ 3.2 Additive Energy Distribution
There are two lines of work that inspire the choice of additive energy distributions. Firstly, these distributions have been used to enhance compositionality in generative tasks  (Du et al., 2020, 2021; Liu et al., 2021) but they have not been used in discriminative compositionality. Secondly, for readers from the causal machine learning community, it may be useful to think of additive energy distributions from the perspective of the independent mechanisms principle (Janzing and Schölkopf, 2010; Parascandolo et al., 2018). The principle states that the data distribution is composed of independent data generation modules, where the notion of independence refers to algorithmic independence and not statistical independence. In these distributions, we think of energy functions of different attributes as independent functions.
@@
Thus $\sigma(z)$ will be a sparse vector of length $md$ containing $m$ ones.
We also define a vector valued map $E(x)=[E_{1}(x,1),\ldots,E_{1}(x,d),\ldots,E_{m}(x,1),\ldots,E_{m}(x,d)]^{\top}$ where $E_{i}(x,z_{i})$ is the energy term for $i^{th}$ attribute taking the value $z_{i}$.
This allows us to reexpress equation 1 using a simple dot product, denoted $\braket{\cdot,\cdot}$:
@@ 4 Provable Compositional Generalization
Our goal is to learn a distribution $\hat{q}(z|x)$ that matches the test distribution $q(z|x)$ and predict the attributes at test time in a Bayes optimal manner. If we successfully learn the distribution $q(z|x)$, then we can also predict the individual attributes $q(z_{i}|x)$, e.g., the bird class in Waterbirds dataset, by marginalizing over the rest of the attributes, e.g., the background in Waterbirds dataset. Observe that $q(z|x)$ differs from the training $p(z|x)$, which can be estimated through standard ERM with cross-entropy loss. Since some attributes $z$ observed at test time are never observed at train time, the distribution learned via ERM assigns a zero probability to these attributes and thus it cannot match the test distribution $q(z|x)$.
In what follows, we first introduce a novel mathematical object termed Discrete Affine Hull over the set of attributes. We then describe a generative approach for classification that requires us to learn $p(x|z)$ including the partition function, which is not practical. Next, we describe a purely discriminative approach that circumvents the issue of learning $\hat{p}(x|z)$ and achieves the same extrapolation guarantees. We present the generative approach as it allows to understand the results more easily. Building generative models based on our theory is out of scope of this work but is an exciting future work.
@@ 4.1 Discrete Affine Hull
We define the discrete affine hull of a set of attribute vectors $\mathcal{A}=\{z^{(1)},\ldots,z^{(k)}\}$ where $z^{(i)}\in\mathcal{Z}$, as:
@@
In other words, the discrete affine hull of $\mathcal{A}$ is the set of all possible attribute vectors whose one-hot encoding is in the (regular) affine hull of the one-hot encodings of the attribute vectors of $\mathcal{A}$. This construct will be used to characterize what new combinations of attributes we can extrapolate to. We now give a simple example to illustrate discrete affine hull.
Let us revisit the Waterbirds dataset. Suppose we observe data from three out of the four groups. In one-hot encoding, we represent WB as $[1,0]$ and LB as $[0,1]$. We represent Water as $[1,0]$ and Land as $[0,1]$. Below we show that the attribute vector WB on L represented as [1 0 0 1]delimited-[]1001[1\;0\;0\;1][ 1 0 0 1 ] can be expressed as an affine combination of the remaining three attribute vectors. Based on this, we can conclude that the discrete affine hull of three one-hot concatenated vectors contains all the four possible one-hot concatenations.
@@
In Section 9.4, we generalize the above finding and develop a mathematical characterization of discrete affine hulls that leads to an easy recipe to visualize these sets. In the remainder whenever we use affine hull it means discrete affine hull.
@@ 4.2 Extrapolation of Conditional Density
If the true and learned distribution ($p(\cdot|z)$ and $\hat{p}(\cdot|z)$) are additive energy distributions, then $\hat{p}(\cdot|z)=p(\cdot|z),\forall z\in\mathcal{Z}^{\mathsf{train}}\implies\hat{p}(\cdot|z^{\prime})=p(\cdot|z^{\prime}),\forall z^{\prime}\in\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$.
The result above argues that so long as the group $z^{\prime}$ is in the discrete affine hull of $\mathcal{Z}^{\mathsf{train}}$, the estimated density extrapolates to it.
@@
$\begin{split}\braket{\sigma(z^{\prime}),\hat{E}(x)}=\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}\Big{(}\braket{\sigma(z),E(x)}+C(z)\Big{)}=\braket{\sigma(z^{\prime}),E(x)}+\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}C(z),\end{split}$ (5)
@@
Using extrapolation of conditional density for compositional generalization of classification.
@@
$q(z^{\prime}|x)=\frac{q(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}q(x|z^{\prime\prime})q(z^{\prime\prime})}=\frac{p(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}p(x|z^{\prime\prime})q(z^{\prime\prime})}\approx\frac{\hat{p}(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}\hat{p}(x|z^{\prime\prime})q(z^{\prime\prime})}$
where we used the property of compositional shifts $q(x|z)=p(x|z)$. If we know test group prior $q(z^{\prime})$ (or e.g. assume it to be uniform), we can directly use the expression in RHS to correctly compute the test group probabilities $q(z|x)$, even those for attribute combinations never seen at training.
@@ 4.3 Extrapolation of Discriminative Model
In Section 4.2, we saw how we could, in principle, obtain a classifier that generalizes under compositional shift, by first training conditional probability density models $\hat{p}(x|z)$. But high dimensional probability density modeling remains very challenging, and involves dealing with intractable partition functions. It is typically deemed much simpler to learn a discriminative classifier.
Can we achieve the same extrapolation without having to estimate the entire distribution of $x$ conditional on $z$? This question brings us to our method, which we refer to as Compositional Risk Minimization (CRM).
@@
$p(z|x)=\frac{p(x|z)p(z)}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}p(x|z^{\prime})p(z^{\prime})}=\frac{\exp\Big{(}-\braket{\sigma(z),E(x)}+\log p(z)-\log\mathbb{Z}(z)\Big{)}}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(z^{\prime}),E(x)}+\log p(z^{\prime})-\log\mathbb{Z}(z^{\prime})\Big{)}}$
@@
$\tilde{p}(z|x)=\frac{\exp\bigg{(}-\braket{\sigma(z),\tilde{E}(x)}+\log\hat{p}(z)-\tilde{B}(z)\bigg{)}}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}\exp\bigg{(}-\braket{\sigma(z^{\prime}),\tilde{E}(x)}+\log\hat{p}(z^{\prime})-\tilde{B}(z^{\prime})\bigg{)}},$ (6)
where $\hat{p}(z)$ is the empirical estimate of the prior over $z$, i.e., $p(z)$, $\tilde{E}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{md}$ is a function to be learned, bias $\tilde{B}$ is a lookup table containing a learnable offset for each combination of attribute. Given a data point $(x,z)$, loss $\ell(z,\tilde{p}(\cdot|x))=-\log\tilde{p}(z|x)$ measures the prediction performance of $\tilde{p}(\cdot|x)$. The risk is defined as the expected loss as follows
$R(\tilde{p})=\mathbb{E}_{(x,z)\sim p}\Big{[}\ell(z,\tilde{p}(\cdot|x))\Big{]}=\mathbb{E}_{(x,z)\sim p}\Big{[}-\log\tilde{p}(z|x)\Big{]}.$ (7)
@@
$\hat{E},\hat{B}\in\operatorname*{arg\,min}_{\tilde{E},\tilde{B}}R(\tilde{p}).$ (8)
If the minimization is over arbitrary functions, then $\hat{p}(\cdot|x)=p(\cdot|x),\forall x\in\mathbb{R}^{n}$. In the second step of CRM, we compute our final predictor $\hat{q}(z|x)$ as follows. Let $\hat{q}(z)$ be an estimate of the marginal distribution over the attributes $q(z)$ with support $\hat{\mathcal{Z}}^{\mathsf{test}}$. We define, for each $z\in\mathcal{Z}^{\mathsf{test}}$,
$\hat{q}(z|x)=\frac{\exp\bigg{(}-\braket{\sigma(z),\hat{E}(x)}+\log\hat{q}(z)-B^{\star}(z)\bigg{)}}{\sum_{z^{\prime}\in\hat{\mathcal{Z}}^{\mathsf{test}}}\exp\bigg{(}-\braket{\sigma(z^{\prime}),\hat{E}(x)}+\log\hat{q}(z^{\prime})-B^{\star}(z^{\prime})\bigg{)}},$ (9)
@@
$B^{\star}(z)=\log\Bigg{(}\mathbb{E}_{x\sim p(x)}\Bigg{[}\frac{\exp\Big{(}-\braket{\sigma(z),\hat{E}(x)}\Big{)}}{\sum_{\tilde{z}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(\tilde{z}),\hat{E}(x)}+\log p(\tilde{z})-\hat{B}(\tilde{z})\Big{)}}\Bigg{]}\Bigg{)}$ (10)
where $\hat{E}$, $\hat{B}$ are the solutions from optimization equation 8. Each of these steps is easy to operationalize. We explain the process and provide pseudocode in Section 5 .

### v2

@@ Abstract
Compositional generalization is a crucial step towards developing data-efficient intelligent machines that generalize in human-like ways. In this work, we tackle a challenging form of distribution shift, termed compositional shift, where some attribute combinations are completely absent at training but present in the test distribution. This shift tests the model’s ability to generalize compositionally to novel attribute combinations in discriminative tasks. We model the data with flexible additive energy distributions, where each energy term represents an attribute, and derive a simple alternative to empirical risk minimization termed compositional risk minimization (CRM). We first train an additive energy classifier to predict the multiple attributes and then adjust this classifier to tackle compositional shifts. We provide an extensive theoretical analysis of CRM, where we show that our proposal extrapolates to special affine hulls of seen attribute combinations. Empirical evaluations on benchmark datasets confirms the improved robustness of CRM compared to other methods from the literature designed to tackle various forms of subpopulation shifts.
@@
We work with multi-attribute data, where each input (e.g., an image) is associated with multiple categorical attributes, and the task is to predict an attribute or multiple attributes. During training, we observe inputs from only a subset of all possible combinations of individual attributes, and during test we will see novel combinations of attributes never seen at training. Following Liu et al. (2023), we refer to this distribution shift as compositional shift. Towards the goal of tackling these compositional shifts, we develop an adaptation of naive discriminative Empirical Risk Minimization (ERM) tailored for multi-attribute data under compositional shifts. We term our approach Compositional Risk Minimization (CRM). The foundations of CRM are built on additive energy distributions that are studied in generative compositionality (Liu et al., 2022a), where each energy term represents one attribute. In CRM, we first train an additive energy classifier to predict all the attributes jointly, and then we adjust this classifier for compositional shifts.
@@
  • Theory of discriminative compositional shifts: For the family of additive energy distributions, we prove that additive energy classifiers generalize compositionally to novel combinations of attributes represented by a special mathematical object, which we call discrete affine hull. Our characterization of extrapolation is sharp, i.e., we show that it is not possible to generalize beyond discrete affine hull. We show that the volume of discrete affine hull grows very fast in the number of training attribute combinations thus generalizing to many attribute combinations. The proof techniques developed in this work are very different from existing works on distribution shifts and hence may be of independent interest.
@@
  • A practical method: CRM is a simple algorithm for training classifiers, which first trains an additive energy classifier and then adjusts the trained classifier for tackling compositional shifts. We empirically validate the superiority of CRM to other methods previously proposed for addressing subpopulation shifts.
@@ 1 Problem Setting
1.1 Generalizing under Compositional Shifts
@@
We denote the support of each attribute component $z_{i}$ under training distribution as $\mathcal{Z}_{i}^{\mathsf{train}}$ and the support of $z$ under training distribution as $\mathcal{Z}^{\mathsf{train}}$. The corresponding supports for the test distribution are denoted as $\mathcal{Z}_{i}^{\mathsf{test}}$ and $\mathcal{Z}^{\mathsf{test}}$. We define the Cartesian product of marginal support under training as ${\mathcal{Z}^{\times}}\coloneq\mathcal{Z}_{1}^{\mathsf{train}}\times\mathcal{Z}_{2}^{\mathsf{train}}\times\cdots\mathcal{Z}_{m}^{\mathsf{train}}$.
In this work, we study compositional shifts from a training distribution $p$ to a test distribution $q$, characterized by:
@@
The first point states that the conditional density of inputs conditioned on attributes remains invariant from train to test, which can be understood as the data generation mechanism from attributes to the inputs remains invariant. What changes between train and test is thus due to only shifting prior probabilities of attributes from $p(z)$ to $q(z)$. The second point specifies how these differ in their support: at test we observe novel combinations of individual attributes but not a completely new individual attribute. The task of compositional generalization is then to build classifiers that are robust to such compositional distribution shifts. Also, we remark that the above notion should remind the reader of the notion of Cartesian Product Extrapolation (CPE) from Lachapelle et al. (2024). Specifically, if a model succeeds on test distributions $q(z)$ with support equal to the full Cartesian product ($\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$), then it is said to achieve CPE.
@@ 1.2 Additive Energy Distribution
Thus $\sigma(z)$ will be a sparse vector of length $md$ containing $m$ ones. We also define a vector valued map $E(x)=[E_{1}(x,1),\ldots,E_{1}(x,d),\ldots,E_{m}(x,1),\ldots,E_{m}(x,d)]^{\top}$ where $E_{i}(x,z_{i})$ is the energy term for $i^{th}$ attribute taking the value $z_{i}$. This allows us to reexpress equation 1 using a simple dot product, denoted $\braket{\cdot,\cdot}$:
@@
There are two lines of work that inspire the choice of additive energy distributions. Firstly, these distributions have been used to enhance compositionality in generative tasks  (Du et al., 2020, 2021; Liu et al., 2021) but they have not been used in discriminative compositionality. Secondly, for readers from the causal machine learning community, it may be useful to think of additive energy distributions from the perspective of the independent mechanisms principle (Janzing and Schölkopf, 2010; Parascandolo et al., 2018). The principle states that the data distribution is composed of independent data generation modules, where the notion of independence refers to algorithmic independence and not statistical independence. In these distributions, we think of energy function of an attribute as an independent function.
This is the right juncture to contrast AEDs with distributional assumptions in recent provable approaches to compositional generalization (Dong and Ma, 2022; Wiedemer et al., 2023, 2024; Brady et al., 2023; Lachapelle et al., 2024). These works assume labeling functions or decoders that are deterministic and additive over individual features, proving generalization over the Cartesian product of feature supports (further discussion in Appendix 7). While insightful, this assumption is restrictive, as each attribute combination corresponds to a single observation with limited generative interactions. In contrast, AEDs capture stochastic decoders, offering a more flexible way to model inputs as a conjunction of characteristics (see Appendix 8).
@@ 2 Provable Compositional Generalization
Our goal is to learn a model yielding a $\hat{q}(z|x)$ that will match the test distribution $q(z|x)$ and thus allow us to predict the attributes at test time in a Bayes optimal manner. If we successfully learn the distribution $q(z|x)$, then we can straightforwardly predict the individual attributes $q(z_{i}|x)$, e.g., the bird class in Waterbirds dataset, by marginalizing over the rest, e.g., the background in Waterbirds dataset. Observe that $q(z|x)$ differs from the training $p(z|x)$, which can be estimated through standard ERM with cross-entropy loss. Since some attributes $z$ observed at test time are never observed at train time, the distribution learned via ERM assigns a zero probability to these attributes and thus it cannot match the test distribution $q(z|x)$.
In what follows, we first introduce a novel mathematical object termed Discrete Affine Hull over the set of attributes. We then describe a generative approach for classification that requires us to learn $p(x|z)$ including the partition function, which is not practical. Next, we describe a purely discriminative approach that circumvents the issue of learning $\hat{p}(x|z)$ and achieves the same extrapolation guarantees as the generative approach. We present the generative approach as it allows to understand the results more easily. Building generative models based on our theory is out of scope of this work but is an exciting future work.
@@ 2.1 Discrete Affine Hull
We define the discrete affine hull of a set of attribute vectors $\mathcal{A}=\{z^{(1)},\ldots,z^{(k)}\}$ where $z^{(i)}\in\mathcal{Z}$ as:
@@
In other words, the discrete affine hull of $\mathcal{A}$ consists of all attribute vectors whose one-hot encoding lies in the (regular) affine hull of the one-hot encodings of the attribute vectors of $\mathcal{A}$. This construct helps characterize which new attribute combinations we can extrapolate to.
As an illustration, consider the Waterbirds dataset, where we observe three of four possible groups. In one-hot encoding, WB is $[1,0]$, LB is $[0,1]$, Water is $[1,0]$, and Land is $[0,1]$. We show that the missing attribute vector WB on L, represented as [1 0 0 1]delimited-[]1001[1\;0\;0\;1][ 1 0 0 1 ], can be expressed as an affine combination of the observed vectors, meaning the discrete affine hull of three one-hot concatenated vectors contains all four possible combinations.
@@
In Section 10.5, we generalize this finding to a formal mathematical characterization of discrete affine hulls, providing a visualization method. In Section 10.7, we show how discrete affine hulls generalize the extrapolation of additive functions studied in Dong and Ma (2022); Lachapelle et al. (2024) over discrete domains. We also show how our results lead to a sharp characterization of extrapolation of these functions. Throughout, “affine hull” refers to the discrete affine hull.
@@ 2.2 Extrapolation of Conditional Density
If the true and learned distribution ($p(\cdot|z)$ and $\hat{p}(\cdot|z)$) are AED, then $\hat{p}(\cdot|z)=p(\cdot|z),\forall z\in\mathcal{Z}^{\mathsf{train}}\implies\hat{p}(\cdot|z^{\prime})=p(\cdot|z^{\prime}),\forall z^{\prime}\in\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$.
The result above argues that so long as the group $z^{\prime}$ is in the discrete affine hull of $\mathcal{Z}^{\mathsf{train}}$, the estimated density extrapolates to it. We provide a proof sketch ahead, with the complete proof in Section 10.1.
@@
$\begin{split}\braket{\sigma(z^{\prime}),\hat{E}(x)}=\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}\Big{(}\braket{\sigma(z),E(x)}+C(z)\Big{)}\\ =\braket{\sigma(z^{\prime}),E(x)}+\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}C(z),\end{split}$ (5)
@@
Classifier based on conditional density $p(x|z)$.
@@
$q(z^{\prime}|x)$ $=\frac{q(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}q(x|z^{\prime\prime})q(z^{\prime\prime})}$ (6)
$=\frac{p(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}p(x|z^{\prime\prime})q(z^{\prime\prime})}\approx\frac{\hat{p}(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}\hat{p}(x|z^{\prime\prime})q(z^{\prime\prime})}$
where we used the property of compositional shifts $q(x|z)=p(x|z)$. If we know test group prior $q(z^{\prime})$ (or e.g. assume it to be uniform), we can directly use the expression in RHS to correctly compute the test group probabilities $q(z|x)$, even for groups never seen at training.
@@ 2.3 Extrapolation of Discriminative Model
[…]
But is it really necessary to precisely model the conditional density of high dimensional $x$, when our goal is simply to predict a few classes and attributes $z$, given $x$? We will now develop an alternative approach, Compositional Risk Minimization (CRM), that achieves a similar extrapolation result as Theorem 1, while being based on simple discriminative classifier training. It sidesteps the need and difficulties of explicitly modeling $p(x|z)$ and doesn’t require dealing with the partition function throughout training.
@@
$p(z|x)=\frac{p(x|z)p(z)}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}p(x|z^{\prime})p(z^{\prime})}$
$=\frac{\exp\Big{(}-\braket{\sigma(z),E(x)}+\log p(z)-\log\mathbb{Z}(z)\Big{)}}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(z^{\prime}),E(x)}+\log p(z^{\prime})-\log\mathbb{Z}(z^{\prime})\Big{)}}$
@@
$\tilde{p}(z|x)=\frac{\exp\Big{(}-\braket{\sigma(z),\tilde{E}(x)}+\log\hat{p}(z)-\tilde{B}(z)\Big{)}}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(z^{\prime}),\tilde{E}(x)}+\log\hat{p}(z^{\prime})-\tilde{B}(z^{\prime})\Big{)}}$ (7)
where $\hat{p}(z)$ is the empirical estimate of the prior over $z$, i.e., $p(z)$, $\tilde{E}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{md}$ is a function to be learned, bias $\tilde{B}$ is a lookup table containing a learnable offset for each combination of attribute. Given a data point $(x,z)$, loss $\ell(z,\tilde{p}(\cdot|x))=-\log\tilde{p}(z|x)$ measures the prediction performance of $\tilde{p}(\cdot|x)$. The risk, defined as the expected loss, corresponds to the negated conditional log-likelihood:
$R(\tilde{p})=\mathbb{E}_{(x,z)\sim p}\big{[}\ell(z,\tilde{p}(\cdot|x))\big{]}=\mathbb{E}_{(x,z)\sim p}\big{[}-\log\tilde{p}(z|x)\big{]}$ (8)
@@
$\hat{E},\hat{B}\in\operatorname*{arg\,min}_{\tilde{E},\tilde{B}}R(\tilde{p})$ (9)
If the minimization is over arbitrary functions, then $\hat{p}(\cdot|x)=p(\cdot|x),\forall x\in\mathbb{R}^{n}$. In the second step of CRM, we compute our final predictor $\hat{q}(z|x)$ as follows. Let $\hat{q}(z)$ be an estimate of the marginal distribution over the attributes $q(z)$ with support $\hat{\mathcal{Z}}^{\mathsf{test}}$. For each $z\in\mathcal{Z}^{\mathsf{test}}$
$\hat{q}(z|x)=\frac{\exp\Big{(}-\braket{\sigma(z),\hat{E}(x)}+\log\hat{q}(z)-B^{\star}(z)\Big{)}}{\sum_{z^{\prime}\in\hat{\mathcal{Z}}^{\mathsf{test}}}\exp\Big{(}-\braket{\sigma(z^{\prime}),\hat{E}(x)}+\log\hat{q}(z^{\prime})-B^{\star}(z^{\prime})\Big{)}}$ (10)
@@
$B^{\star}(z)=\log\mathbb{E}_{x\sim p}\Big{[}\frac{\exp\Big{(}-\braket{\sigma(z),\hat{E}(x)}\Big{)}}{\sum_{\tilde{z}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(\tilde{z}),\hat{E}(x)}+\log p(\tilde{z})-\hat{B}(\tilde{z})\Big{)}}\Big{]}$ (11)
where $\hat{E}$, $\hat{B}$ are the solutions from optimization equation 9. Note that $\hat{B}(z)$ was learned for all $z\in\mathcal{Z}^{\mathrm{train}}$ but never for $z\in\mathcal{Z}^{\mathrm{test}}$, hence the necessity of extrapolation $B^{*}$. Each of these steps is easy to operationalize. We explain the process and provide pseudocode in Section 3.